# FedHK-MVFC: Federated Heat Kernel Multi-View Clustering

### v1

@@ 6.4.2 Ablation Studies
The analysis reveals that both clients achieve perfect clustering performance, with no misclassifications observed. This highlights the effectiveness of the federated learning setup in leveraging diverse data distributions across clients. The visualization of Hospital A clustering results are shown in Figure 7, where each client’s clustering performance is represented by a separate confusion matrix. The matrices indicate that both clients successfully identify the four phenotypes, with high precision and recall across all categories.
@@
The visualization of Hospital B clustering results are shown in Figure 8. The clustering results indicate that both clients successfully identify the four phenotypes, with high precision and recall across all categories.
@@ 6.4.3 Communication Efficiency Analysis
For federated experiments, we analyze communication costs:
Metric FedMV-KMeans FedHK-MVFC Reduction
@@
Figure 10: Communication Efficiency Comparison
The substantial reduction in communication overhead while maintaining faster convergence demonstrates the efficiency of our federated heat kernel-enhanced approach.
@@ 6.4.4 Broader Implications and Future Extensions
Our synthetic multi-view data generation framework contributes to the research community by providing:
  • Reproducible Benchmarks: Standardized synthetic datasets with known ground truth for fair algorithm comparison
  • Scalable Generation: Framework easily adaptable to different numbers of views, clusters, and sample sizes
  • Geometric Diversity: Comprehensive coverage of shape complexities relevant to real-world applications
  • Federated Scenarios: Built-in support for federated learning experimental setups
@@
The geometric patterns in our synthetic framework mirror structures found in real applications:
  • Medical Imaging: Organ boundaries (crescent-like), vascular networks (spiral), lesions (elliptical)
  • Social Networks: Community structures (ring-like), influence patterns (radial)
  • Genomics: Gene expression clusters (complex non-convex shapes)
  • Computer Vision: Object boundaries with varying curvatures and topologies
Algorithmic Insights:
The experimental results on synthetic data reveal several key algorithmic insights:
@@
    Heat Kernel Effectiveness: The superior performance on non-convex shapes (crescent, heart) demonstrates that heat kernel distances capture intrinsic geometric properties that Euclidean distances miss.
@@
    Multi-View Synergy: The complementary view design shows that effective multi-view clustering requires careful balance between view correlation and information redundancy.
@@
    Federated Robustness: Minimal performance degradation in federated settings indicates that our approach maintains geometric sensitivity even with distributed data.
@@
    Scalability Properties: Linear scaling with dataset size and sub-linear communication complexity make the approach practical for large-scale applications.
Framework Extensions:
The synthetic data generation framework can be extended to support:
Multi-scale geometries: $\mathcal{G}_{multi}=\bigcup_{s\in\mathcal{S}}s\cdot\mathcal{G}_{base}$ (137)
Dynamic clusters: $\mathcal{G}_{t}=\mathcal{T}_{t}(\mathcal{G}_{t-1})$ (138)
Noise heterogeneity: $\sigma_{i,h}=f(\mathbf{x}_{i},h,\text{context})$ (139)
These extensions enable evaluation of algorithm performance under more complex scenarios including multi-scale data, temporal evolution, and adaptive noise patterns.
7 Conclusion
This work has introduced a comprehensive framework for heat kernel-enhanced multi-view clustering in both centralized and federated settings. Through the development of sophisticated synthetic data generation methodologies and rigorous experimental evaluation, we have demonstrated several key contributions:
Theoretical Contributions:
  • Novel integration of quantum field theory concepts (heat-kernel coefficients) with practical clustering algorithms
  • Mathematically rigorous framework for multi-view data generation with controlled geometric complexity
  • Theoretical analysis of federated clustering with privacy guarantees and communication efficiency
Algorithmic Innovations:
  • HK-MVFC algorithm with closed-form update rules and proven convergence properties
  • FedHK-MVFC federated extension with minimal performance degradation (98.2% retention)
  • Heat kernel distance transformation that captures intrinsic geometric structures
Experimental Validation:
  • Comprehensive synthetic data framework with eight distinct geometric patterns across two views
  • Substantial performance improvements (8-12% accuracy gain) over state-of-the-art baselines
  • 70% reduction in communication overhead while maintaining clustering quality
  • Rigorous statistical validation and ablation studies confirming design choices
Future research directions include extension to dynamic multi-view scenarios, integration with deep learning architectures, and application to large-scale real-world datasets in healthcare, social networks, and computer vision. The heat kernel-enhanced framework provides a solid theoretical foundation for these extensions while maintaining computational efficiency and privacy preservation requirements.
@@ 7 Conclusion
Our work demonstrates that principled integration of mathematical concepts from quantum field theory with practical machine learning algorithms can yield substantial performance improvements while maintaining theoretical rigor. This approach opens new avenues for algorithm development that bridge fundamental mathematics with contemporary distributed computing challenges.

### v2

@@ 6.6.2 Ablation Studies
The analysis reveals that both clients achieve perfect clustering performance, with no misclassifications observed. This highlights the effectiveness of the federated learning setup in leveraging diverse data distributions across clients. The visualization of Hospital A clustering results is shown in Figure 9, where each client’s clustering performance is represented by a separate confusion matrix. The matrices indicate that both clients successfully identify the four phenotypes, with high precision and recall across all categories.
@@
The visualization of Hospital B clustering results is shown in Figure 10. The clustering results indicate that both clients successfully identify the four phenotypes, with high precision and recall across all categories.
@@ 6.6.3 Communication Efficiency Analysis
In the context of federated experiments, an analysis of communication costs is conducted, as outlined in Table 9. As demonstrated, the substantial reduction in communication overhead while maintaining faster convergence demonstrates the efficiency of our federated heat kernel-enhanced approach. The reduction of bytes per round and total communication round is up to 70% as compared to Fed-MVFCM. Convergence rates exhibited a decline, with a 28.1% decrease observed. The findings indicated that FedHK-MVFC exhibited superior performance in comparison to its predecessor with regard to communication and convergence with H-KC-based QFT in the objective function.
Metric Fed-MVFCM FedHK-MVFC Reduction
@@
Table 9: Communication Efficiency Comparison
@@ 6.6.4 Broader Implications and Future Extensions
Our synthetic multi-view data generation framework contributes to the research community by providing reproducible benchmarks through standardized synthetic datasets with known ground truth for fair algorithm comparison. The framework offers scalable generation capabilities that are easily adaptable to different numbers of views, clusters, and sample sizes. It encompasses geometric diversity through comprehensive coverage of shape complexities relevant to real-world applications, while also incorporating built-in support for federated learning experimental setups to facilitate research in distributed multi-view clustering scenarios.
@@
The geometric patterns in our synthetic framework mirror structures found in real applications. In medical imaging, we observe organ boundaries that exhibit crescent-like shapes, vascular networks displaying spiral patterns, and lesions presenting elliptical configurations. Social networks demonstrate community structures with ring-like formations and influence patterns that follow radial distributions. Genomic data reveal gene expression clusters characterized by complex non-convex shapes. Computer vision applications encounter object boundaries with varying curvatures and topologies. These diverse geometric manifestations across different domains validate the practical relevance of our synthetic data generation framework and underscore the necessity for clustering algorithms capable of handling such geometric complexity.
Algorithmic Insights and Framework Extensions:
The experimental results on synthetic data reveal several key algorithmic insights. The superior performance on non-convex shapes (crescent, heart) demonstrates that heat kernel distances capture intrinsic geometric properties that Euclidean distances miss. The complementary view design demonstrates that effective multi-view clustering requires a careful balance between view correlation and information redundancy. Minimal performance degradation in federated settings indicates that our approach maintains geometric sensitivity even with distributed data. Linear scaling with dataset size and sub-linear communication complexity make the approach practical for large-scale applications. Furthermore, the synthetic data generation framework can be extended to support multi-scale geometries $\mathcal{G}_{multi}=\bigcup_{s\in\mathcal{S}}s\cdot\mathcal{G}_{base}$, dynamic clusters $\mathcal{G}_{t}=\mathcal{T}_{t}(\mathcal{G}_{t-1})$, and noise heterogeneity $\sigma_{i,h}=f(\mathbf{x}_{i},h,\text{context})$. These extensions enable the evaluation of algorithm performance under more complex scenarios, including multi-scale data, temporal evolution, and adaptive noise patterns.
7 Conclusion and Future Directions
This work has introduced a comprehensive framework for heat kernel-enhanced multi-view clustering in both centralized and federated settings. Through the development of sophisticated synthetic data generation methodologies and rigorous experimental evaluation, several key contributions have been demonstrated. Our theoretical contributions include the novel integration of quantum field theory concepts (heat-kernel coefficients) with practical clustering algorithms, a mathematically rigorous framework for multi-view data generation with controlled geometric complexity, and theoretical analysis of federated clustering with privacy guarantees and communication efficiency. From an algorithmic perspective, the HK-MVFC algorithm was introduced, complete with closed-form update rules and proven convergence properties. The FedHK-MVFC federated extension was also presented, demonstrating minimal performance degradation (100 percentage point retention). Additionally, a heat kernel distance transformation was proposed to capture intrinsic geometric structures. The experimental validation encompasses a synthetic data framework with eight distinct geometric patterns across two views. The experimental results show substantial performance improvements (1-50 percentage point accuracy gain) over state-of-the-art baselines, a 70 percentage point reduction in communication overhead while maintaining clustering quality, and rigorous statistical validation and ablation studies confirming design choices.
7.1 Limitations
Despite the promising results, several limitations warrant acknowledgment:
@@ 7.1 Limitations
    Parameter Sensitivity: The clustering efficacy is contingent upon the proper tuning of the fuzziness parameter $m$ and the view weight exponent $\alpha$. While our experiments demonstrate robustness across a reasonable parameter range, automated parameter selection mechanisms would enhance practical applicability.
@@
    Computational Complexity: Although our approach achieves linear scaling with dataset size, the heat kernel coefficient computation introduces additional computational overhead compared to conventional Euclidean distance methods. For extremely large-scale applications, further optimization strategies may be necessary.
@@
    Synthetic Data Validation: While our synthetic datasets exhibit controlled geometric complexity and statistical properties, validation on real-world medical datasets is essential to fully assess the framework’s practical utility. The synthetic scenarios, though representative, may not capture all nuances of actual clinical data distributions.
@@
    View Heterogeneity Assumptions: The current framework assumes that all participating clients possess the same set of views with consistent feature spaces. Extensions to handle scenarios where different clients have access to different view subsets would broaden the applicability of the approach.
7.2 Future Research Directions
Building upon the theoretical and experimental foundations established in this work, we identify several promising directions for future investigation:
  1. 1.
    Real-World Medical Data Applications: A critical next step involves the application of FedHK-MVFC to large-scale real-world medical datasets, particularly The Cancer Genome Atlas (TCGA) data repository. TCGA provides multi-omics data (genomic, transcriptomic, proteomic, and clinical) across thousands of cancer patients, presenting an ideal testbed for evaluating heat kernel-enhanced federated clustering in authentic healthcare scenarios with inherent data heterogeneity and privacy constraints.
  2. 2.
    Automated Hyperparameter Optimization: Development of principled strategies for automatic selection of the fuzziness parameter $m$ and view weight exponent $\alpha$ through techniques such as Bayesian optimization, meta-learning, or cross-validation frameworks adapted for federated settings would significantly enhance the framework’s practical usability.
  3. 3.
    Dynamic Multi-View Scenarios: Extension of the framework to handle temporal evolution of multi-view data, where both the number of views and their statistical properties change over time. This would enable applications in longitudinal patient monitoring and time-series clinical data analysis.
  4. 4.
    Integration with Deep Learning Architectures: Exploration of hybrid approaches that combine heat kernel-enhanced clustering with deep representation learning, potentially leveraging graph neural networks or variational autoencoders to learn optimal feature representations before applying federated clustering.
  5. 5.
    Scalability Enhancements: Investigation of distributed computing strategies, including GPU acceleration and asynchronous federated learning protocols, to scale the framework to datasets with millions of samples and hundreds of features across numerous distributed clients.
  6. 6.
    Heterogeneous View Handling: Development of extensions that accommodate scenarios where different clients possess different subsets of views, enabling more flexible federated collaboration in real-world settings where data collection protocols vary across institutions.
  7. 7.
    Privacy-Utility Trade-off Analysis: Comprehensive theoretical and empirical analysis of the privacy-utility trade-offs under varying differential privacy budgets, secure aggregation protocols, and communication constraints to provide principled guidance for privacy-sensitive applications.
@@ 7.2 Future Research Directions
Our work demonstrates that the principled integration of mathematical concepts from quantum field theory with practical machine learning algorithms can yield substantial performance improvements while maintaining theoretical rigor. This approach opens new avenues for algorithm development that bridge fundamental mathematics with contemporary distributed computing challenges. The heat kernel-enhanced framework provides a solid theoretical foundation for the proposed future extensions while maintaining computational efficiency and privacy preservation requirements essential for real-world deployment in sensitive domains such as healthcare, finance, and social networks.
7.3 Extensions to Future Internet and Clustered Federated Learning
The proposed FedHK-MVFC framework naturally extends to emerging Future Internet architectures, where distributed intelligence, edge computing, and decentralized network topologies create novel opportunities and challenges for federated learning. We envision several critical extensions that align with Future Internet paradigms:
7.3.1 Hierarchical Clustered Federated Learning for Network Topologies
Future Internet infrastructures operate across multi-tiered network hierarchies—edge devices, fog nodes, regional data centers, and cloud servers. Our framework can be extended to support hierarchical clustered federated learning where:
$\mathcal{H}_{FedHK}$ $=\{\mathcal{L}_{\text{edge}},\mathcal{L}_{\text{fog}},\mathcal{L}_{\text{regional}},\mathcal{L}_{\text{cloud}}\}$ (136)
$\mathbf{A}_{global}^{(t)}$ $=\text{HierAgg}\left(\{\mathbf{A}_{\ell}^{(t)}\}_{\ell\in\mathcal{H}_{FedHK}}\right)$ (137)
where $\text{HierAgg}(\cdot)$ represents a hierarchical aggregation strategy that respects network topology constraints, bandwidth limitations, and latency requirements. This extension enables clustered federated learning at multiple network scales, where edge clusters aggregate locally before communicating with higher-tier nodes, significantly reducing communication overhead in bandwidth-constrained network environments.
7.3.2 Network-Aware Heat Kernel Coefficients
In Future Internet scenarios, data exhibits network-induced geometric structures reflecting communication patterns, routing topologies, and distributed sensor configurations. We propose network-aware heat kernel coefficients:
$\delta_{[\ell]ij}^{h,\text{net}}=f\left(\delta_{[\ell]ij}^{h},\mathcal{G}_{\text{network}},\mathbf{p}_{latency}\right)$ (138)
where $\mathcal{G}_{\text{network}}$ represents the underlying network graph topology and $\mathbf{p}_{latency}$ captures communication delay patterns. This network-aware formulation enables heat kernel methods to account for the physical communication infrastructure, optimizing clustering strategies based on both data similarity and network proximity.
7.3.3 Interdisciplinary Applications and Cross-Domain Collaboration
The versatility of the FedHK-MVFC framework extends beyond traditional healthcare applications, enabling interdisciplinary collaboration across diverse Future Internet domains:
Smart City Infrastructure: Federated clustering of heterogeneous sensor networks (traffic cameras, environmental monitors, energy grids) enables city-wide pattern recognition while respecting jurisdictional data boundaries and privacy regulations.
Industrial IoT and Manufacturing: Multi-facility collaborative quality control through federated clustering of production line data, equipment sensor readings, and supply chain logistics enables predictive maintenance and process optimization without exposing proprietary manufacturing data.
Financial Services: Cross-institutional fraud detection and risk assessment through federated clustering of transaction patterns, customer behavior, and market indicators while maintaining regulatory compliance and competitive confidentiality.
Environmental Monitoring: Global climate pattern recognition through federated analysis of distributed meteorological stations, satellite imagery, and oceanic sensor networks, enabling collaborative scientific discovery while respecting international data sovereignty.
Social Computing and Network Science: Privacy-preserving community detection in federated social networks, enabling collaborative understanding of social dynamics, information diffusion, and behavioral patterns across multiple platforms without centralizing sensitive user data.
7.3.4 Adaptive Clustered Federated Learning Protocols
Future Internet environments exhibit dynamic characteristics—client availability fluctuates, network conditions vary, and data distributions evolve over time. We propose adaptive protocols that adjust clustering strategies based on network state:
$\mathcal{P}_{adaptive}^{(t)}=\arg\min_{\mathcal{P}}\left\{\mathcal{L}_{clustering}(\mathcal{P})+\lambda_{net}\mathcal{C}_{communication}(\mathcal{P},\mathcal{N}^{(t)})\right\}$ (139)
where $\mathcal{N}^{(t)}$ represents the network state at round $t$, $\mathcal{C}_{communication}$ quantifies communication costs, and $\lambda_{net}$ balances clustering quality against network efficiency. This formulation enables the framework to dynamically adapt to network congestion, client dropouts, and bandwidth variations characteristic of Future Internet environments.
7.3.5 Cross-Domain Transfer Learning for Federated Clustering
Extending beyond single-domain applications, we envision cross-domain transfer mechanisms where knowledge learned in one federated clustering task (e.g., medical imaging) can be transferred to related tasks in different domains (e.g., industrial defect detection) through heat kernel geometry preservation:
$\bm{\delta}_{target}=\mathcal{T}\left(\bm{\delta}_{source},\mathcal{M}_{domain}\right)$ (140)
where $\mathcal{T}(\cdot)$ represents a domain adaptation function and $\mathcal{M}_{domain}$ captures domain-specific characteristics. This capability facilitates interdisciplinary knowledge transfer, enabling federated learning systems to leverage insights across diverse application domains.
These extensions position FedHK-MVFC as a foundational framework for clustered federated learning in Future Internet architectures, supporting interdisciplinary collaboration, network-aware optimization, and adaptive learning protocols essential for next-generation distributed intelligence systems. The integration of heat kernel methods with emerging network paradigms provides a mathematically principled approach to handling the geometric complexity, privacy requirements, and communication constraints inherent in Future Internet environments.
Acknowledgment: The author would like to express her gratitude to Ishtiaq Hussain for conducting the simulations for Co-FKM, MinMax-FCM, WV-Co-FCM, and Co-FW-MVFCM, as presented in Tables 2-7. Additionally, Hussain provided a draft paragraph that outlines the corresponding experimental analysis.
Data availability: Data will be made available on https://github.com/KristinaP09/FedHK-MVFC.
Funding: This research received no external funding.
Conflicts of interest: The author declare no conflict of interest.